# OpenXAI: Towards a Transparent Evaluation of Model Explanations

### v1

@@ Appendix C Synthetic Dataset
Our proposed data generation process described in Algorithm 1 is designed to encapsulate arbitrary feature dependencies, unambiguous definitions of local neighborhoods, and clear descriptions of feature influences. In our algorithm 1, the local neighborhood of a sample is controlled by its cluster membership $k$, while the degree of feature dependency can be easily controlled by the user setting $\Sigma_{k}$ to desired values. The default value of $\Sigma_{k}=\mathbf{I}$, where $\mathbf{I}$ is the identity matrix, indicating that all features are independent of one another. The elements of the true underlying weight vector $w_{k}$ are sampled from a uniform distribution: $w_{k}\sim\mathcal{U}(l,u)$. We set l=1l=-1 and $u=1$. The masking vectors $m_{k}$ with elements $m_{k,j}$ are generated by a Bernoulli distribution: $m_{k,j}\sim\mathcal{B}(p)$, where the parameter $p$ controls the ground-truth explanation sparsity. We set $p=0.25$. Further, the cluster centers are chosen as follows: when the number of features $d$ is smaller than or equal to the number of clusters $K$, then we set the first cluster center to $1,0,\dots,0]$, the second one to $0,1,\dots,0]$, etc. We also introduce a multiplier $\kappa$ to control the distance between the cluster centers so that the cluster centers are located at $\mu_{1}=\kappa\cdot[1,0,\dots,0]$, $\mu_{2}=\kappa\cdot[0,1,\dots,0]$, etc. We set $\kappa=6$. In general, we have observed that lower values of $\kappa$ result in more complex classification problems. When the number of features $d$ is greater than the number of clusters $K$, we adjust the above described approach slightly: we first compute $\ell=\lfloor K/d\rfloor$, and then we compute the cluster centers for the first $d$ clusters as described above, for the next $d$ clusters we use $\mu_{d+1}=2\kappa\cdot[1,0,\dots,0]$, $\mu_{d+2}=2\kappa\cdot[0,1,\dots,0]$, etc. If $\ell=1$, we stop this procedure here, else we continue, and repeat filling up the clusters $\mu_{2d+1}=3\kappa\cdot[1,0,\dots,0]$, $\mu_{2d+2}=3\kappa\cdot[0,1,\dots,0]$, etc.
input : number of clusters: $K$, cluster centers: $\mu_{1},\mu_{2},\dots,\mu_{K}]$, cluster variances: $\Sigma_{1},\dots,\Sigma_{K}]$, Weight vectors: $w_{1},\dots,w_{K}]$, masking vectors: $m_{1},\dots,m_{K}]$
@@
[…]
@@
Unique ground truth explanation for each local neighborhood: The dataset $D$ is said to comprise of unique ground truth explanations for each local neighborhood if the ground truth labels of instances in each cluster $k\in\{1,2\cdots K\}$ are generated as a function of a subset of features $A_{k}\subseteq A$ where there is a clear relative ordering among features in $A_{k}$ which is captured by the weight vector $w_{k}$, and features in $A$ are completely independent of each other. [Note that the influential feature set $A_{k}$ corresponding to the cluster $k$ is also captured using the mask vector $m_{k}$ (See Section 2) where $m_{k,i}=1$ if the $i^{th}$ feature is in set $A_{k}$]. This would imply that there is a clear description of the top-$T$ features (and their relative ordering) where $T\leq|A_{k}|$ that influence the ground truth labeling process.
@@
Appendix D Extending OpenXAI
Figure 5: A snapshot of the OpenXAI leaderboard submission form. Users can use this single form to incorporate new datasets, pre-trained models, explanation methods, and evaluation metrics into OpenXAI framework.
In this section, we walk through how users can extend and contribute to OpenXAI by adding custom datasets, pre-trained models, explanation methods, or evaluation metrics. In Figure 5, we show a snapshot of our FORM that can be used by users to submit requests for new datasets, pre-trained models, explanation methods, or evaluation metrics.
Custom Datasets. Adding new datasets into OpenXAI ’s pipeline is as simple as uploading a .csv file or a .zip folder. Users can submit requests to incorporate new datasets into OpenXAI by providing either the complete dataset (with or without data splits) or individual training and testing files for the respective dataset.
Custom Predictive Models. Users can also submit requests to include new pre-trained models into the OpenXAI framework. Here, users have to provide i) model architecture and implementation details, ii) model hyperparameters and training details, and iii) trained weights for reproducing model performance.
Custom Evaluation Metrics. Users can include new evaluation metrics into OpenXAI by filling out the submit request form and providing the GitHub link to their code and a summary of their metrics. Once approved, the metric should be implemented as a function in the Evaluator class. The input data, ground-truth labels, model predictions, explanations, and the underlying black-box model are all provided in the Evaluator class. The evaluation metric should return a scalar score as a part of their respective implementation.
Custom Explanation Methods. Users can include new explanation methods into OpenXAI ’s list of explainers by following the template provided in Figure 1. The abstract method defined in Figure 1 ensures that the output of new/existing explanation methods returns explanations as a tensor. Note that all custom explanation methods should extend the Explainer class. Finally, users must provide the GitHub link to their code and a summary of their explanation method.
We invite submissions to any one or multiple benchmarks in OpenXAI. To be included in the leaderboard, please fill out this FORM, include results of your explanation method and provide the full implementation of your algorithm with proper documentations and GitHub link.
Appendix E Benchmarking Analysis
E.1 Real-World Datasets
@@
In addition to synthetic dataset, OpenXAI library includes five real-world benchmark datasets from high-stakes domains. Table 1 provides a summary of the real-world datasets currently included in OpenXAI. Our library implements multiple data split strategies and allows users to customize the percentages of train-test splits. To this end, if a given dataset comes with a pre-determined train and test splits, OpenXAI loads the training and testing dataloaders from those pre-determined splits. Otherwise, OpenXAI’s dataloader divides the entire dataset randomly into train (70%) and test (30%). Next, we detail the covariates $\mathbf{x}$ and labels $\mathbf{y}$ for each dataset.
@@
German Credit. The dataset comprises of demographic (age, gender), personal (marital status), and financial (income, credit duration) features from 1,000 credit applicants, where they are categorized into good vs. bad customer depending on their credit risk [14].
COMPAS. The dataset has criminal records and demographics features for 18,876 defendants who got released on bail at the U.S state courts during 1990-2009. The task is to classify defendants into bail (i.e., unlikely to commit a violent crime if released) vs. no bail (i.e., likely to commit a violent crime) [27].
Adult Income. The dataset contains demographic (e.g., age, race, and gender), education (degree), employment (occupation, hours-per week), personal (marital status, relationship), and financial (capital gain/loss) features for 48,842 individuals. The task is to predict whether an individual’s income exceeds \$50K per year vs. not [66].
Give Me Some Credit. The dataset incorporates demographic (age), personal (number of dependents), and financial (e.g., monthly income, debt ratio, etc.) features for 250,000 individuals. The task is to predict the probability that a customer will experience financial distress vs. not in the next two years. The aim of the dataset is to build models that customers can use to the best financial decisions [18].
@@
Home Equity Line of Credit (HELOC). The dataset comprises of financial (e.g., total number of trades, average credit months in file) attributes from anonymized applications submitted by 9,871 real homeowners. A HELOC is a line of credit typically offered by a bank as a percentage of home equity. The fundamental task is to use the information about the applicant in their credit report to predict whether they will repay their HELOC account within 2 years [17].
E.2 Explanation Methods
@@
OpenXAI provides implementations for six explanation methods: Vanilla Gradients [56], Integrated Gradients [60], SmoothGrad [59], Gradient x Input [55], LIME [51], and SHAP [44]. Table E.2 summarizes how these methods differ along two axes: whether each method requires “White-box Access” to model gradients and whether each method learns a local approximation model to generate explanations.
@@
Implementations. We used existing public implementations of all explanation methods in our experiments. We used the following captum software package classes: i) captum.attr.Saliency for Vanilla Gradients; ii) captum.attr.IntegratedGradients for Integrated Gradients; iii) captum.attr.NoiseTunnel and captum.attr.Saliency for SmoothGrad; iv) captum.attr.InputXGradient for Gradient x Input; and v) captum.attr.KernelShap for SHAP. Finally, we use the authors’ LIME python package for LIME.
E.3 Hyperparameter details
OpenXAI uses default hyperparameter settings for all explanation methods following the authors’ guidelines. Every explanation method has a corresponding parameter dictionary that stores all the default parameter values. Below, we detail the hyperparameters of individual explanation methods used in our experiments.
a) params_lime = {‘lime_mode’: ‘tabular’, ‘lime_sample_around_instance’: True, ‘lime_kernel_ width’: 0.75, ‘lime_n_samples’: 1000, ‘lime_discretize_continuous’: False, ‘lime_standard_ deviation’: float(np.sqrt(0.05))}
@@ E.3 Hyperparameter details
b) params_shap = {‘shap_subset_size’: 50}
c) params_grads = {‘grad_absolute_value’: False}
@@
d) params_sg = {‘sg_n_samples’: 500, ‘sg_standard_deviation’: float(np.sqrt(0.05))}
Description. The parameter ‘sg_n_samples’ sets the number of samples used in the Gaussian random variable to smooth the gradient, while ‘sg_standard_deviation’ determines the size of the local neighboorhood the gaussian random variables are sampled from. We use a small value, ensuring that we in fact sample from a local neighborhood.
e) params_ig = {‘ig_method’: ‘gausslegendre’, ‘ig_multiply_by_inputs’: False, ‘ig_baseline’: ‘mean’}
Further, we parameterized our data generating process as follows.
@@
Description. The above parameters can be matched with the parameters from the data generating process described in Appendix C: in particular, ‘sparsity’=p=p, ‘upper_weight’ =u=u, ‘lower_weight’ =l=l, ‘dim’=d=d, ‘n_clusters’ =K=K, ‘distance_to_center’ $\kappa$ and ‘sigma’=None implies that the identity matrix is used, i.e., $\Sigma_{k}=\mathbf{I}$.
@@
Finally, we denote the top-$k$ value using {‘percentage_most_important’: 0.25}, i.e., we consistently use 25% of the top features for calculating our metric scores.
E.4 Model details
Our current release of OpenXAI has two pre-trained models: i) a logistic regression model and ii) a deep neural network model for all datasets. To support systematic, reproducible, and efficient benchmarking of post hoc explanation methods, we provide the model weights for both models trained on all six datasets in our pipeline. For neural network models, we use two fully connected hidden layers with 100 nodes in each layer, with ReLU activation functions and an output softmax layer for all datasets. We train both models for 50 epochs using an Adam optimizer with a learning rate of $\eta{=}0.001$. Next, we show the model performance of both models on all six datasets.
Table 7: Results of the machine learning models trained on six datasets. Shown are the accuracy of LR and ANN models trained the datasets. The best performance is bolded.
@@ E.4 Model details
83.0%
72.0%
72.0%
85.4%
84.0%
94.1%
92.0%
75.0%
74.0%
85.0%
85.0%
95.0%
E.5 Additional Results on LR models
Table 8: Ground-truth and predicted faithfulness results on the SynthGauss dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@ E.5 Additional Results on LR models
[…]
Table 9: Ground-truth and predicted faithfulness results on the German credit dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
Table 10: Ground-truth and predicted faithfulness results on the COMPAS dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across all instances in the test set. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.

### v5

@@ Appendix C Synthetic Dataset
Our proposed data generation process described in Algorithm 1 is designed to encapsulate arbitrary feature dependencies, unambiguous definitions of local neighborhoods, and clear descriptions of feature influences. In our algorithm 1, the local neighborhood of a sample is controlled by its cluster membership $k$, while the degree of feature dependency can be easily controlled by the user setting $\Sigma_{k}$ to desired values. The default value of $\Sigma_{k}=\mathbf{I}$, where $\mathbf{I}$ is the identity matrix, indicating that all features are independent of one another. The elements of the true underlying weight vector $w_{k}$ are sampled from a uniform distribution: $w_{k}\sim\mathcal{U}(l,u)$. We set $l=-1$ and $u=1$. The masking vectors $m_{k}$ with elements $m_{k,j}$ are generated by a Bernoulli distribution: $m_{k,j}\sim\mathcal{B}(p)$, where the parameter $p$ controls the ground-truth explanation sparsity. We set $p=0.25$. Further, the cluster centers are chosen as follows: when the number of features $d$ is smaller than or equal to the number of clusters $K$, then we set the first cluster center to $[1,0,\dots,0]$, the second one to $[0,1,\dots,0]$, etc. We also introduce a multiplier $\kappa$ to control the distance between the cluster centers so that the cluster centers are located at $\mu_{1}=\kappa\cdot[1,0,\dots,0]$, $\mu_{2}=\kappa\cdot[0,1,\dots,0]$, etc. We set $\kappa=6$. In general, we have observed that lower values of $\kappa$ result in more complex classification problems. When the number of features $d$ is greater than the number of clusters $K$, we adjust the above described approach slightly: we first compute $\ell=\lfloor K/d\rfloor$, and then we compute the cluster centers for the first $d$ clusters as described above, for the next $d$ clusters we use $\mu_{d+1}=2\kappa\cdot[1,0,\dots,0]$, $\mu_{d+2}=2\kappa\cdot[0,1,\dots,0]$, etc. If $\ell=1$, we stop this procedure here, else we continue, and repeat filling up the clusters $\mu_{2d+1}=3\kappa\cdot[1,0,\dots,0]$, $\mu_{2d+2}=3\kappa\cdot[0,1,\dots,0]$, etc.
input :  number of clusters: $K$, cluster centers: $[\mu_{1},\mu_{2},\dots,\mu_{K}]$, cluster variances: $[\Sigma_{1},\dots,\Sigma_{K}]$, Weight vectors: $[w_{1},\dots,w_{K}]$, masking vectors: $[m_{1},\dots,m_{K}]$
@@
       $y_{i}=\mathbb{I}(\Pi[i]>\tilde{\pi})$     # Make sure classes are balanced ;
@@
Unique ground truth explanation for each local neighborhood: The dataset $D$ is said to comprise of unique ground truth explanations for each local neighborhood if the ground truth labels of instances in each cluster $k\in\{1,2\cdots K\}$ are generated as a function of a subset of features $A_{k}\subseteq A$ where there is a clear relative ordering among features in $A_{k}$ which is captured by the weight vector $w_{k}$, and features in $A$ are completely independent of each other. [Note that the influential feature set $A_{k}$ corresponding to the cluster $k$ is also captured using the mask vector $m_{k}$ (See Section 2) where $m_{k,i}=1$ if the $i^{th}$ feature is in set $A_{k}$]. This would imply that there is a clear description of the top-$T$ features (and their relative ordering) where $T\leq|A_{k}|$ influence the ground truth labeling process.
@@
Appendix D Benchmarking Analysis
D.1 Real-World Datasets
@@
In addition to the synthetic dataset, the OpenXAI library includes seven real-world benchmark datasets from high-stakes domains. Table 1 provides a summary of the real-world datasets currently included in OpenXAI. Our library implements multiple data split strategies and allows users to customize the percentages of train-test splits. To this end, if a given dataset comes with a pre-determined train and test splits, OpenXAI loads the training and testing dataloaders from those pre-determined splits. Otherwise, OpenXAI’s dataloader divides the entire dataset randomly into train (80%) and test (20%). Next, we detail the covariates $\mathbf{x}$ and labels $\mathbf{y}$ for each dataset.
@@
German Credit [Dua and Graff, 2017] The dataset comprises of demographic (age, gender), personal (marital status), and financial (income, credit duration) features from 1,000 credit applicants, where they are categorized into good vs. bad customer depending on their credit risk.
@@
HELOC [FICO, 2022] The dataset comprises of financial (e.g., total number of trades, average credit months in file) attributes from anonymized applications submitted by 9,871 real homeowners. A HELOC (Home Equity Line of Credit) is a line of credit typically offered by a bank as a percentage of home equity. The fundamental task is to use the information about the applicant in their credit report to predict whether they will repay their HELOC account within 2 years.
Adult Income [Yeh and Lien, 2009] The dataset contains demographic (e.g., age, race, and gender), education (degree), employment (occupation, hours-per week), personal (marital status, relationship), and financial (capital gain/loss) features for 45,222 individuals. The task is to predict whether an individual’s income exceeds \$50K per year vs. not.
COMPAS [Jordan and Freiburger, 2015] The dataset has criminal records and demographics features for 6,172 defendants released on bail at U.S state courts during 1990-2009. The task is to classify defendants into bail (i.e., unlikely to commit a violent crime if released) vs. no bail (i.e., likely to commit a violent crime).
Give Me Some Credit [Freshcorn, 2022] The dataset incorporates demographic (age), personal (number of dependents), and financial (e.g., monthly income, debt ratio, etc.) features for 102,209 individuals. The task is to predict the probability whether or not a customer will experience financial distress in the next two years. The aim of the dataset is to build models that customers can use to the best financial decisions.
We also include the Pima-Indians Diabetes [Smith et al., 1988] and Framingham Heart Study [Fra, ] datasets (see Appendix D.5).
D.2 Explanation Methods
@@
OpenXAI provides implementations for six explanation methods: Vanilla Gradients [Simonyan et al., 2014], Integrated Gradients [Sundararajan et al., 2017], SmoothGrad [Smilkov et al., 2017], Gradient x Input [Shrikumar et al., 2017], LIME [Ribeiro et al., 2016], and SHAP [Lundberg and Lee, 2017b]. Table D.2 summarizes how these methods differ along two axes: whether each method requires “White-box Access” to model gradients and whether each method learns a local approximation model to generate explanations.
@@
Implementations. We used existing public implementations of all explanation methods in our experiments. We used the following captum software package classes: i) captum.attr.Saliency for Vanilla Gradients; ii) captum.attr.IntegratedGradients for Integrated Gradients; iii) captum.attr.NoiseTunnel and captum.attr.Saliency for SmoothGrad; iv) captum.attr.InputXGradient for Gradient x Input; and v) captum.attr.KernelShap for SHAP. Finally, we use the authors’ python package for LIME.
D.3 Hyperparameter details
OpenXAI uses default hyperparameter settings for all explanation methods following the authors’ guidelines. Every explanation method has a corresponding parameter dictionary that stores all the default parameter values. Below, we detail the hyperparameters of individual explanation methods used in our experiments. Seed is set to 0 for all explanation methods to allow reproducibility.
a) params_lime = {‘seed’: 0, ‘lime_mode’: ‘tabular’, ‘sample_around_instance’: True, ‘kernel_ width’: 0.75, ‘n_samples’: 1000, ‘discretize_continuous’: False, ‘std’: 0.1}
@@ D.3 Hyperparameter details
b) params_shap = {‘seed’: 0, ‘n_samples’: 500, ‘model_impl’: ‘torch’}
Description. The parameter ‘shap_n_samples’ controls the number of samples of the original model used to train the surrogate SHAP model, while the parameter ‘model_impl’ toggles between models from ‘torch’ and ‘sklearn’.
c) params_grads = {‘absolute_value’: False}
@@
d) params_sg = {‘n_samples’: 500, ‘standard_deviation’: 0.1}
Description. The parameter ‘sg_n_samples’ sets the number of samples used in the Gaussian random variable to smooth the gradient, while ‘sg_standard_deviation’ determines the size of the local neighborhood the Gaussian random variables are sampled from. We use a small value, ensuring that we sample from a local neighborhood.
e) params_ig = {‘method’: ‘gausslegendre’, ‘multiply_by_inputs’: False}
Further, we parameterized our data-generating process as follows.
@@
Description. The above parameters can be matched with the parameters from the data generating process described in Appendix C: in particular, ‘sparsity’$=p$, ‘upper_weight’ $=u$, ‘lower_weight’ $=l$, ‘$=d$, ‘n_clusters’ $=K$, ‘distance_to_center’ $=\kappa$ and ‘sigma’=None implies that the identity matrix is used, i.e., $\Sigma_{k}=\mathbf{I}$.
@@
Finally, we dynamically infer the top-$k$ value from the input parameter $k$ and the total number of features, $d$, as follows: top-$k=k$ for $\{k\in\mathbb{Z}\mid 1\leq k\leq d\}$; top-$k=d$ for $k=-1$; and top-$k=\lceil k\times d\rceil$ for $\{k\in\mathbb{R}\mid 0<k<1\}$.
For predictive faithfulness, we set $\sigma{=}0.1$, and sample 100 perturbations. The flip percentage is set to $\sigma\sqrt{2/\pi}$ in all cases, which leads to the same expected magnitude of change between continuous and discrete variables when drawing from a Normal distribution. For stability, we set $\sigma{=}10^{-5}$, and sample 1000 points before selecting a subset of 100 samples that share the original input’s prediction.
The $p$ norm for all stability metrics is set to 2. Seed is set to the special case of -1 for our evaluation metrics, which assumes the index of the specific test instance being evaluated as the seed being used. All default hyperparameters used in our experiments can also be found in the experiment_config.json file of our OpenXAI GitHub repository.
D.4 Model details
Our current release of OpenXAI has two pre-trained models: i) a logistic regression model and ii) a deep neural network model for all datasets. To support systematic, reproducible, and efficient benchmarking of post hoc explanation methods, we provide the model weights for both models trained on all eight datasets in our pipeline. For neural network models, we use two fully connected hidden layers with 100 nodes in each layer, with ReLU activation functions and an output softmax layer for all datasets. We train both models for 100 epochs using an Adam optimizer with a learning rate of $\eta{=}0.001$. We use min-max scaling to normalize train/test splits, along with dataset-specific batch sizes/positive class weights (to balance the loss function). The best epoch in terms of test accuracy is selected, excluding an initial warmup phase of 5 epochs, and provided that dataset-specific bounds are not exceeded for the proportion of the majority class predicted. Next, we show the performance of both models on all eight datasets.
Table 7: Results of the machine learning models trained on eight datasets. Shown are the test accuracies of LR and ANN models. Values corresponding to best performance are bolded.
@@ D.4 Model details
Pima-Indians Diabetes
Framingham Heart Study
81.76%
64.00%
72.35%
85.26%
83.25%
93.79%
69.48%
80.19%
90.32%
61.50%
73.92%
85.02%
84.62%
93.43%
77.92%
80.87%
D.5 Results for Pima-Indians Diabetes and Framingham Heart Study datasets
We include two new datasets from the healthcare domain and benchmark state-of-the-art explanation methods on these datasets. More specifically, we include the Pima-Indians Diabetes [Smith et al., 1988] and Framingham Heart Study [Fra, ] datasets both of which have been utilized in recent XAI research. New results with these datasets are shown below. We observe similar insights with these new datasets: predictive faithfulness results are shown in Tables  8-D.5 below); stability results are shown in Tables D.5-D.5 below.
Table 8: Ground-truth and predictive faithfulness results on the Pima Indians Diabetes dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@ D.5 Results for Pima-Indians Diabetes and Framingham Heart Study datasets
0.470$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.492$\pm$0.01
1.000$\pm$0.00
0.488$\pm$0.01
0.980$\pm$0.00
-0.076$\pm$0.03
1.000$\pm$0.00
1.000$\pm$0.00
-0.047$\pm$0.03
1.000$\pm$0.00
-0.047$\pm$0.03
0.986$\pm$0.00
0.154$\pm$0.02
1.000$\pm$0.00
1.000$\pm$0.00
0.195$\pm$0.02
1.000$\pm$0.00
0.195$\pm$0.02
0.985$\pm$0.00
0.109$\pm$0.02
1.000$\pm$0.00
1.000$\pm$0.00
0.133$\pm$0.02
1.000$\pm$0.00
0.135$\pm$0.02
0.985$\pm$0.00
0.075$\pm$0.02
1.000$\pm$0.00
1.000$\pm$0.00
0.195$\pm$0.02
1.000$\pm$0.00
0.195$\pm$0.02
0.985$\pm$0.00
0.054$\pm$0.01
1.000$\pm$0.00
1.000$\pm$0.00
0.133$\pm$0.02
1.000$\pm$0.00
0.135$\pm$0.02
0.985$\pm$0.00
0.031$\pm$0.00
0.026$\pm$0.00
0.026$\pm$0.00
0.030$\pm$0.00
0.026$\pm$0.00
0.030$\pm$0.00
0.026$\pm$0.00
0.013$\pm$0.00
0.021$\pm$0.00
0.021$\pm$0.00
0.015$\pm$0.00
0.021$\pm$0.00
0.015$\pm$0.00
0.021$\pm$0.00
Table 9: Ground-truth and predictive faithfulness results on the Framingham Heart dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
@@
Table 10: Predictive faithfulness results on the Pima Indians Diabetes dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.084$\pm$0.008
0.055$\pm$0.006
0.055$\pm$0.006
0.058$\pm$0.006
0.054$\pm$0.006
0.069$\pm$0.007
0.054$\pm$0.006
0.034$\pm$0.004
0.074$\pm$0.007
0.074$\pm$0.007
0.070$\pm$0.007
0.074$\pm$0.007
0.056$\pm$0.006
0.074$\pm$0.007
Table 11: Predictive faithfulness results on the Framingham Heart dataset for all explanation methods with ANN model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method PGU ($\downarrow$) PGI ($\uparrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
0.130$\pm$0.005
0.110$\pm$0.005
0.112$\pm$0.005
0.118$\pm$0.005
0.110$\pm$0.005
0.121$\pm$0.005
0.110$\pm$0.005
0.047$\pm$0.003
0.089$\pm$0.004
0.085$\pm$0.004
0.079$\pm$0.004
0.088$\pm$0.004
0.073$\pm$0.004
0.088$\pm$0.004
Table 12: Stability results on the Pima Indians Diabetes dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.86$\pm$0.02
-1.01$\pm$0.06
-2.12$\pm$0.06
0.70$\pm$0.02
6.81$\pm$0.06
6.72$\pm$0.04
7.17$\pm$0.02
13.11$\pm$0.04
-2.04$\pm$0.08
-3.20$\pm$0.08
2.75$\pm$0.10
7.52$\pm$0.08
7.77$\pm$0.08
8.37$\pm$0.05
Table 13: Stability results on the Framingham Heart dataset for all explanation methods with LR model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.76$\pm$0.01
1.25$\pm$0.02
0.43$\pm$0.01
1.41$\pm$0.01
8.59$\pm$0.01
9.64$\pm$0.01
8.95$\pm$0.01
15.47$\pm$0.02
0.80$\pm$0.04
-0.30$\pm$0.02
3.55$\pm$0.08
11.22$\pm$0.03
11.73$\pm$0.02
10.67$\pm$0.02
Table 14: Stability results on the Pima Indians Diabetes dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.86$\pm$0.02
4.37$\pm$0.70
1.95$\pm$0.19
4.59$\pm$0.76
10.85$\pm$0.09
10.31$\pm$0.04
10.83$\pm$0.11
11.90$\pm$0.01
4.54$\pm$0.76
1.93$\pm$0.22
4.77$\pm$0.80
10.93$\pm$0.08
10.36$\pm$0.04
11.02$\pm$0.12
15.06$\pm$0.10
8.42$\pm$0.99
2.91$\pm$0.15
8.72$\pm$0.99
13.83$\pm$0.10
13.36$\pm$0.16
13.94$\pm$0.15
Table 15: Stability results on the Framingham Heart dataset for all explanation methods with ANN model. Shown are log average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.
Method RIS ($\downarrow$) RRS ($\downarrow$) ROS ($\downarrow$)
Random
VanillaGrad
IntegratedGrad
Gradient x Input
SmoothGrad
SHAP
LIME
11.76$\pm$0.01
5.06$\pm$0.17
4.11$\pm$0.17
5.09$\pm$0.18
26.14$\pm$0.30
16.77$\pm$0.32
26.56$\pm$0.48
11.25$\pm$0.01
4.19$\pm$0.16
3.23$\pm$0.21
4.21$\pm$0.17
25.94$\pm$0.29
16.41$\pm$0.31
26.52$\pm$0.47
15.21$\pm$0.05
4.50$\pm$0.37
5.34$\pm$0.33
4.87$\pm$0.29
26.77$\pm$0.31
19.49$\pm$0.29
27.43$\pm$0.47
D.6 Remaining Results on LR models
Table 16: Ground-truth and predictive faithfulness results on the Synthetic dataset for all explanation methods with LR model. Shown are average and standard error metric values computed across 1000 test instances. $\uparrow$ indicates that higher values are better, and $\downarrow$ indicates that lower values are better. Values corresponding to best performance are bolded.